# Escalating LLM-based Code Translation Benchmarking into the Class-level Era

### v2

@@ Abstract.
To achieve this, we construct a class-level code translation benchmark, ClassEval-T, and make the first attempt to extensively assess recent LLMs’ performance on class-level code translation. ClassEval-T is extended from ClassEval, a well-known class-level Python code generation benchmark consisting of multiple practical coding topics, such as database operation and game design, and diverse contextual dependencies (e.g., fields, methods, and libraries). It cost us 360 person-hours to accomplish the manual migration to Java and C++ with complete code samples and associated test suites. Subsequently, we design three translation strategies (i.e., holistic, min-dependency, and standalone) for class-level code translations and evaluate six recent LLMs of hybrid, code, and general kinds in diverse families and sizes on ClassEval-T. Experimental results demonstrate a remarkable performance drop compared with the most widely studied method-level code translation benchmark, and obvious discrepancies among LLMs appear, showing the effectiveness of ClassEval-T in measuring recent LLMs. Afterwards, we further discuss the usage scenarios for diverse translation strategies and LLMs’ ability to dependency awareness when translating class samples. Finally, 1,397 failure cases made by the best-performing LLM under test are thoroughly analyzed and categorized in this paper for practical guidance and future enlightenment.
@@ 1. Introduction
Automated code translation aims to migrate codebases from one Programming Language (PL) to another to satisfy the diverse extensions (Liu et al., 2023; Nguyen et al., 2015; Wu et al., 2010), which is critically important for coding productivity. In recent years, with the advancement of Large Language Models (LLMs), the correctness and readability of code translation have achieved substantial improvement (Yang et al., 2024a; Yuan et al., 2024; Nitin and Ray, 2024). According to a series of recent studies (Yang et al., 2024a; Pan et al., 2024), LLMs, such as CodeLlama (Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023) and GPT-3.5 (OpenAI, 2024), can correctly translate over 70%-80% code samples with the most basic prompt among diverse translation pairs, even for those domain-specific scenarios (Tao et al., 2024), such as Python-to-Scala. However, previous studies mainly experiment with statement/method-level code translation benchmarks, consisting of shorter code lengths, limited dependencies, and only algorithmic problems. Therefore, they cannot reflect the code translation capability of LLMs in practical development, and it is highly motivated and necessary to construct a brand-new code translation benchmark aligning with coding practice for LLMs’ assessment, thereby providing insightful research direction and practical guidance for academia and industry.
@@
Empirical Study: Based on ClassEval-T, we make the first attempt to extensively assess recent LLMs’ performance on class-level code translation. Specifically, six recent LLMs of diverse kinds (e.g., hybrid, code, and general), families (e.g., CodeLlama (Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023), Llama (Llama Team, 2024), and Gemma (Team et al., 2024)), and sizes (ranging from 7B to 236B) are involved. For each studied LLM, we evaluate their performance with three distinctive translation strategies, i.e., holistic translation (translating the whole class at once), min-dependency translation (translating each module of a class one by one, given the minimal necessary dependencies), and standalone translation (translating each module of a class one by one without dependency information). For each translated code sample, we evaluate its correctness via Computational Accuracy (CA) and Compilation Success Rate (CSR). Besides, we also incorporate DEPendency evaluation metrics (DEP) to investigate LLMs’ awareness of necessary context during the translation for class-level samples and thoroughly analyze and categorize 1,397 Failed cases at the end of this experiment.
Main findings: Based on our results, we have the following findings. (1) All LLMs perform dramatically worse on class-level code translations than method-level ones, where the latter cannot even tell different LLMs’ performance discrepancies. (2) DeepSeek-V2.5 exhibits predominately superior performance on class-level code translation, while smaller LLMs normally perform worse, given other factors the same. Besides, All LLMs perform better on Python-oriented translations, and Code LLMs do not necessarily outperform general LLMs. (3) DeepSeek-V2.5 always works better on the holistic translation strategy, while the selection of translation strategy for smaller LLMs depends on different scenarios. In addition, Python-oriented translations work better on the holistic strategy, while C++-oriented translations prefer the min-dependency strategy. (4) The holistic translation strategy improves LLMs’ awareness of field dependencies, while the min-dependency strategy allows LLMs to invoke more necessary libraries. As for method dependencies, the above two strategies perform neck-to-neck. (5) Syntax errors remain a primary issue in class-level translations, but class-related errors, such as function/variable usage and consistency issues, are more prominent in C++/Java-oriented translations. Based on the above findings, several implications are summarized for both researchers and practitioners. The contributions of this work can be three-fold:
@@
(3) We summarize a series of findings and manually analyze 1,397 failed cases for categorization, shedding light on practical guidance and future research directions.
@@ 2.1. Large Language Models for Code Translation
Large Language Models (LLMs), with more than billions of parameters trained on general textual/code corpora and instructions, have achieved state-of-the-art performance in many coding tasks (Xue et al., 2024a; Ahmed and Devanbu, 2022; Xue et al., 2024b; Li et al., 2024c, e, a; Jiang et al., 2024; Li et al., 2024b). In recent years, owing to the great demand for codebase migrations (Yang et al., 2024b; Tao et al., 2024), automated code translation with LLMs has also attracted increasingly substantial attention. For example, Pan et al. (Pan et al., 2024) investigated the performance of five LLMs in code translation tasks and identified and summarized syntax and semantic errors in their translations. Eniser et al. (Eniser et al., 2024) presented a comprehensive study on LLM-based translation to Rust, where five LLMs are evaluated. Additionally, other studies have explored ways to enhance LLM performance in code translation tasks. For example, Yang et al. (Yang et al., 2024a) introduced UniTrans, which utilizes test cases to improve LLM translation performance. Yuan et al. (Yuan et al., 2024) proposed a novel LLM-based multi-agent system called TRANSAGENT, which enhances LLM-based code translation by correcting syntax and semantic errors through the synergy of four LLM-based agents. Pan et al. (Pan et al., 2023) introduced SteloCoder, a decoder-only StarCoder-based LLM designed specifically for Python-oriented translation. Yin et al. (Yin et al., 2024) proposed a general corrector, Rectifier, a lightweight and universal model for repairing translation errors by learning from mistakes made by existing LLMs. However, despite these studies demonstrating the promise of using LLMs for code translation, their evaluations have primarily been conducted on statement- or method-level benchmarks, leaving a notable lack of studies on code translations for more challenging scenarios. Considering most practical coding is object-oriented (Du et al., 2024), this work constructs a class-level benchmark, ClassEval-T, with longer code lengths, diverse dependencies, and practical coding problems for investigation.
@@ 2.2. Existing Benchmarks for Code Translation
Code translation plays a critical role in cross-language codebase migration. In recent years, a series of benchmarks were successively constructed, typically providing parallel corpora of diverse PLs with accompanied test suites for correctness validation. For example, Puri et al. (Puri et al., [n. d.]) built CodeNet, which covers widely used PLs and includes code translation corpus collected from programming contest sites with extensive metadata. AVATAR (Ahmad et al., 2023) is a collection of 9,515 programming problems with solutions written in Java and Python, sourced from competitive programming sites, online platforms, and open-source repositories. TransCoder-test (Roziere et al., 2020) is the most extensively assessed code translation test set consisting of 948 parallel samples in C++, Java, and Python, derived from GeeksForGeeks(Geeks for Geeks, 2024). However, it only contains 568 samples with test suites in at least one PL. Owing to its huge noise, Yang et al. (Yang et al., 2024a) cleaned this dataset and released a new version. Building on existing datasets, Jiao et al. (Jiao et al., 2023) constructed G-TransEval by extracting parallel functions from program-level parallel code, covering five PLs across four different types. Pan et al. (Pan et al., 2024) manually constructed equivalent Java test suites via JUnit(JUnit 5, 2024) from the Python test suites implemented in EvalPlus (Liu et al., 2024d), a code synthesis evaluation framework extended from HumanEval (Chen et al., 2021), for the validation of Python-to-Java translation. Besides, Tao et al. (Tao et al., 2024), focusing on more diverse and domain-specific PLs, constructed PolyHumanEval, which extends HumanEval into a multilingual benchmark across 14 PLs. Table 1 provides an overview of the above benchmarks, including their release time, parallel PLs, whether they align with practical development, granularity, benchmark size (#Tasks), number of methods in total (#Methods), the average number of test cases per task (#Tests/T), code scale (#LOC/T: average lines of code per task, #Tokens/T: average number of tokens per task, split by blanks), and the number(ratio) of diverse dependencies (#FD: Field dependencies, #MD: Method dependencies, #LD: Library dependencies) across different benchmarks. We also present our constructed benchmark ClassEval-T in the last row for comparison.
@@
Based on Table 1, we have the following observations. First, existing benchmarks typically focus on statement- or method-level code translation tasks (Column “Granularity”). Thus, such tasks often involve a limited number of lines (17.9 on average) and tokens (86.57 on average), which may not fully explore the capacity of recent LLMs that can process longer sequences. Second, statement- or method-level code units lack field and method dependencies (Column “#FD” and “#MD”) or provide minimal library dependencies (Column “#LD”). Because they are normally crawled from online judge websites, such as AtCoder (AtCoder, 2024) and GeeksForGeeks (Geeks for Geeks, 2024), leading them to mainly focus on translating standalone code units without consideration of other code contexts. However, in practical scenarios, methods often depend on each other or share variables. Previous study (Yu et al., 2024) indicated that in open-source projects, only about 30% of methods operate independently of other code contexts. Therefore, the capability of LLMs to translate class-level code involving interdependent methods remains unclear.
@@
Benchmark Time Parallel Programming Languages Practical Granularity #Tasks #Methods #Tests/T #LOC/T #Tokens/T #FD #MD #LD
@@
G-TransEval (Jiao et al., 2023) 2023 C#,C++,JavaScript,Java,Python N Method-level 400 400 5 14.3 93.6 - - -
EvalPlus (Liu et al., 2024d) 2023 Python $\rightarrow$ Java N Method-level 164 164 16.1 9.3 30.6 - - 7(4.3%)
@@ 3.2. The superiority of ClassEval-T
ClassEval-T features a lot of superiority over its previous counterparts, where its detailed statistics are shown in the last row of Table 1. (1) It consists of 94 different practical coding tasks manually designed by Du et al. (Du et al., 2024), including the student registration system, area calculator, gomoku game, etc. (2) As an evaluation benchmark for code translation, ClassEval-T contains more lines (66.7) and tokens (199.5) per task, outperforming previous benchmarks on average by 272.63% and 130.46% respectively. (3) Besides, ClassEval-T carries more diverse dependencies to fields, methods, and libraries. As shown in Table 1, 65.3% of the code samples in ClassEval-T contain field dependencies, 25.9% of them contain method dependencies, and 19.2% of them contain library dependencies. (4) To ensure the effectiveness of validation, ClassEval-T comprises 33.8 test cases per task, outperforming previous benchmarks by 229.76%. At the same time, it also carries extremely high coverage scores on statements (99.7%) and branches (98.2%). Consequently, ClassEval-T demonstrates a much higher complexity and comprehensive test suites that can be used for the assessment of state-of-the-art LLMs.

### v3

@@ Abstract.
To achieve this, we construct a class-level code translation benchmark, ClassEval-T, and make the first attempt to extensively assess recent LLMs’ performance on class-level code translation. ClassEval-T is extended from ClassEval, a well-known class-level Python code generation benchmark consisting of multiple practical coding topics, such as database operation and game design, and diverse contextual dependencies (e.g., fields, methods, and libraries). It cost us 360 person-hours to accomplish the manual migration to Java and C++ with complete code samples and associated test suites. Subsequently, we design three translation strategies (i.e., holistic, min-dependency, and standalone) for class-level code translations and evaluate eight recent LLMs of commercial, general, and code kinds in diverse families and sizes on ClassEval-T. Experimental results demonstrate a remarkable performance drop compared with the most widely studied method-level code translation benchmark, and obvious discrepancies among LLMs appear, showing the effectiveness of ClassEval-T in measuring recent LLMs. Afterwards, we further discuss the usage scenarios for diverse translation strategies and LLMs’ ability to dependency awareness when translating class samples. Finally, 1,243 failure cases made by the best-performing LLM under test are thoroughly analyzed and categorized in this paper for practical guidance and future enlightenment.
@@ 1. Introduction
Automated code translation seeks to efficiently migrate codebases between Programming Languages (PLs) to meet the diverse needs of various platforms, such as desktop applications, websites, and mobile apps. This is essential for enhancing coding productivity and facilitating the extension of business platforms (Liu et al., 2023; Nguyen et al., 2015; Wu et al., 2010). As software development progresses, the need to port code from one PL to another has become increasingly important. In recent years, with the advancement of Large Language Models (LLMs), the correctness and readability of code translation have achieved substantial improvement (Yang et al., 2024a; Yuan et al., 2024; Nitin and Ray, 2024). According to a series of recent studies (Yang et al., 2024a; Pan et al., 2024), LLMs, such as CodeLlama (Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023) and GPT-3.5 (OpenAI, 2024), can correctly translate over 70%-80% of code samples with the most basic prompt among diverse translation pairs, even for those domain-specific scenarios (Tao et al., 2024), such as Python-to-Scala. However, previous studies in code translation primarily experiment with statement/method-level code translation benchmarks, which typically involve shorter code snippets, limited dependencies, and a focus on algorithmic problems. These benchmarks, while useful, fail to capture the complexities encountered in real-world software development. In practice, code translation tasks often involve longer code structures, more intricate dependencies, and diverse functionalities. Therefore, the current benchmarks do not adequately reflect the code translation capabilities of LLMs in real-world coding scenarios. This gap highlights the urgent need for a brand-new code translation benchmark aligning with coding practice for LLMs’ assessment, thereby providing insightful research direction and practical guidance for academia and industry.
@@
Empirical Study: Based on ClassEval-T, we make the first attempt to extensively assess recent LLMs’ performance on class-level code translation. Specifically, eight recent LLMs of diverse kinds (e.g., commercial, general, and code) and sizes (ranging from 7B to 671B) are involved. For each studied LLM, we evaluate their performance with three distinctive translation strategies, i.e., holistic translation (translating the whole class at once), min-dependency translation (translating each module of a class one by one, given the minimal necessary dependencies), and standalone translation (translating each module of a class one by one without dependency information). For each translated code sample, we evaluate its correctness via Computational Accuracy (CA) and Compilation Success Rate (CSR). Besides, we also incorporate DEPendency evaluation metrics (DEP) to investigate LLMs’ awareness of necessary context during the translation for class-level samples and thoroughly analyze and categorize 1,243 failed cases at the end of this experiment.
Main findings: Based on our results, we have the following findings. (1) All LLMs perform dramatically worse on class-level code translations than method-level ones, where the latter cannot even tell different LLMs’ performance discrepancies. (2) Commercial LLMs (e.g, DeepSeek-V3, GPT-4o, and Claude-3.5-Sonnet) exhibit predominately superior performance on class-level code translation, while smaller LLMs normally perform worse, given other factors the same. Besides, All LLMs perform better on Python-oriented translations, and code LLMs do not necessarily outperform general LLMs. (3) Commercial LLMs always work better on the holistic translation strategy, while the selection of translation strategy for smaller LLMs depends on different scenarios. Specifically, as for Python- and Java-oriented translations, smaller LLMs work better on the holistic strategy, while for C++-oriented translations, they perform neck-to-neck on holistic and min-dependency strategies. (4) The holistic translation strategy improves LLMs’ awareness of field dependencies, while the min-dependency strategy allows LLMs to invoke more necessary libraries. As for method dependencies, the above two strategies perform neck-to-neck. (5) Syntax errors remain a primary issue in class-level translations, but class-related errors, such as function/variable usage and consistency issues, are more prominent in C++/Java-oriented translations. Based on the above findings, several implications are summarized for both researchers and practitioners. The contributions of this work can be three-fold:
@@
(3) We summarize a series of findings and manually analyze 1,243 failed cases for categorization, shedding light on practical guidance and future research directions.
@@ 2.1. Large Language Models for Code Translation
Large Language Models (LLMs), with more than billions of parameters trained on general textual/code corpora and instructions, have achieved state-of-the-art performance in many coding tasks (Xue et al., 2024a; Ahmed and Devanbu, 2022; Xue et al., 2024b; Li et al., 2024c, e, a; Jiang et al., 2024; Li et al., 2024b). In recent years, owing to the great demand for codebase migrations (Yang et al., 2024b; Tao et al., 2024), automated code translation with LLMs has also attracted increasingly substantial attention. For example, Pan et al. (Pan et al., 2024) investigated the performance of five LLMs in code translation tasks and identified and summarized syntax and semantic errors in their translations. Eniser et al. (Eniser et al., 2024) presented a comprehensive study on LLM-based translation to Rust, where five LLMs are evaluated. Additionally, other studies have explored ways to enhance LLM performance in code translation tasks. For example, Yang et al. (Yang et al., 2024a) introduced UniTrans, which utilizes test cases to improve LLM translation performance. Yuan et al. (Yuan et al., 2024) proposed a novel LLM-based multi-agent system called TRANSAGENT, which enhances LLM-based code translation by correcting syntax and semantic errors through the synergy of four LLM-based agents. Pan et al. (Pan et al., 2023) introduced SteloCoder, a decoder-only StarCoder-based (Li et al., 2023) LLM designed specifically for Python-oriented translation. Yin et al. (Yin et al., 2024) proposed a general corrector, Rectifier, a lightweight and universal model for repairing translation errors by learning from mistakes made by existing LLMs. However, despite these studies demonstrating the promise of using LLMs for code translation, their evaluations have primarily been conducted on statement- or method-level benchmarks, leaving a notable lack of studies on code translations for more challenging scenarios. Considering most practical coding is object-oriented (Du et al., 2024), this work constructs a class-level benchmark, ClassEval-T, with longer code lengths, diverse dependencies, and practical coding problems for investigation.
@@ 2.2. Existing Benchmarks for Code Translation
Code translation plays a critical role in cross-language codebase migration. In recent years, a series of benchmarks were successively constructed, typically providing parallel corpora of diverse PLs with accompanied test suites for correctness validation. For example, Puri et al. (Puri et al., [n. d.]) built CodeNet, which covers widely used PLs and includes code translation corpus collected from programming contest sites with extensive metadata. AVATAR (Ahmad et al., 2023) is a collection of programming problems with solutions written in Java and Python, sourced from competitive programming sites, online platforms, and open-source repositories. TransCoder-test (Roziere et al., 2020) is the most extensively assessed code translation test set consisting of 948 parallel samples in C++, Java, and Python, derived from GeeksForGeeks(Geeks for Geeks, 2024). However, it only contains 568 samples with test suites in at least one PL. Owing to its huge noise, Yang et al. (Yang et al., 2024a) cleaned this dataset and released a new version. Building on existing datasets, Jiao et al. (Jiao et al., 2023) constructed G-TransEval by extracting parallel functions from program-level parallel code, covering five PLs across four different types. Yan et al. (Yan et al., 2023b) constructed CodeTransOcean, a large-scale comprehensive benchmark, where the MultilingualTrans dataset supports translation between eight popular PLs. However, even though the MultilingualTrans test set contains over 7,000 samples, after our filtering process, only 44 tasks simultaneously cover all parallel PLs and primarily focus on the translation of basic syntactic structures to algorithmic implementations. xCodeEval (Khan et al., 2024) is an execution-based multilingual multitask evaluation benchmarks, derived from data collected from Codeforces (Cod, [n. d.]), and can be utilized for tasks such as code generation and code translation. Besides, Tao et al. (Tao et al., 2024), focusing on more diverse and domain-specific PLs, constructed PolyHumanEval, which extends HumanEval into a multilingual benchmark across 14 PLs. Table 1 provides an overview of the above benchmarks, including their release time, parallel PLs, whether they align with practical development, granularity, benchmark size (#Tasks), average number of methods per PL (#Methods/P), the average number of test cases per task (#Tests/T), code scale (#LOC/T: average lines of code per task, #Tokens/T: average number of tokens per task, split by blanks), and the number(ratio) of diverse dependencies (#FD: Field dependencies, #MD: Method dependencies, #LD: Library dependencies) across different benchmarks. We also present our constructed benchmark ClassEval-T in the last row for comparison.
Additionally, there are some benchmarks for code translation tasks that do not include parallel PLs. For example, Pan et al. (Pan et al., 2024) manually constructed equivalent Java test suites via JUnit(JUnit 5, 2024) from the Python test suites implemented in EvalPlus (Liu et al., 2024d), a code synthesis evaluation framework extended from HumanEval (Chen et al., 2021), for the validation of Python-to-Java translation. The NicheTrans test set in CodeTransOcean (Yan et al., 2023b) focuses on translations from domain-specific PLs to popular PLs. Yan et al. (Yan et al., 2023a) introduced CodeScope, an execution-based, multilingual, multitask evaluation benchmark crawled from Codeforces for measuring LLM capabilities on coding tasks including code translation. Furthermore, there are also two concurrent works focusing on repository-level code translation. Ibrahimzada et al. (Ibrahimzada et al., 2024) proposed AlphaTrans to translate ten real-world open-source projects from Java to Python. Ou et al. (Ou et al., 2024) introduced a repository-level code translation benchmark comprising 375 tasks targeting Rust. However, owing to the absence of parallel PLs in these benchmarks, two main limitations appear. Firstly, researchers cannot fairly compare models’ translation performance among different PLs, as samples of different PLs implementing different functionalities. Secondly, parallel samples can be considered as an assurance that samples can be translated from one PL to another. Without parallel samples, untranslatable samples may present in the test set. Once models fail in certain cases, we cannot identify whether it is owing to models’ weaknesses or the untranslatability of cases, thereby inducing biased and unreliable evaluations. As a result, we consider the above benchmarks to be flawed for evaluation, and we only include and report benchmarks that contain parallel PLs for comparison in Table 1.
@@
Based on Table 1, we have the following observations. First, existing benchmarks typically focus on statement- or method-level code translation tasks (Column “Granularity”). Thus, such tasks often involve a limited number of lines (21.8 on average) and tokens (95.2 on average), which may not fully explore the capacity of recent LLMs that can process longer sequences. Second, statement- or method-level code units lack field and method dependencies (Column “#FD” and “#MD”) or provide minimal library dependencies (Column “#LD”). Because they are normally crawled from online judge websites, such as AtCoder (Inc., [n. d.]), Codeforces (Cod, [n. d.]), and GeeksForGeeks (Geeks for Geeks, 2024), leading them to mainly focus on translating standalone code units without consideration of other code contexts. However, in practical scenarios, methods often depend on each other or share variables. Previous study (Yu et al., 2024) indicated that in open-source projects, only about 30% of methods operate independently of other code contexts. Therefore, the capability of LLMs to translate class-level code involving interdependent methods remains unclear.
@@
Benchmark Time Parallel Programming Languages Practical Granularity #Tasks #Methods/P #Tests/T #LOC/T #Tokens/T #FD #MD #LD
@@
G-TransEval (Jiao et al., 2023) 2023 C++,C#,Java,Python,JavaScript N Method-level 400 400 5 14.3 93.6 - - -
CodeTransOcean-
MultilingualTrans (Yan et al., 2023b)
2023
C,C++,C#,Java,Python,
Go,PHP,Visual Basic
N Statement/Method-level 44 57.9 1 23.8 82.1 - 8(27.6%) 3(10.3%)
xCodeEval-test (Khan et al., 2024) 2024
C,C++,C#,Java,Python,Ruby,Go,
JavaScript,Kotlin,PHP,Rust
N Statement/Method-level 226 406.1 46.8 30.8 95.5 - 24(20.2%) 27(20.1%)
@@ 3.2. The superiority of ClassEval-T
ClassEval-T features a lot of superiority over its previous counterparts, where its detailed statistics are shown in the last row of Table 1. (1) It consists of 94 different practical coding tasks manually designed by Du et al. (Du et al., 2024), including the student registration system, area calculator, gomoku game, etc. (2) As an evaluation benchmark for code translation, ClassEval-T contains more lines (66.7) and tokens (199.5) per task, outperforming previous benchmarks on average by 205.96% and 109.56% respectively. (3) Besides, ClassEval-T carries more diverse dependencies to fields, methods, and libraries. As shown in Table 1, 65.3% of the code samples in ClassEval-T contain field dependencies, 25.9% of them contain method dependencies, and 19.2% of them contain library dependencies. (4) To ensure the effectiveness of validation, ClassEval-T comprises 33.8 test cases per task, outperforming previous benchmarks by 154.14%. At the same time, it also carries extremely high coverage scores on statements (99.7%) and branches (98.2%). Consequently, ClassEval-T demonstrates a much higher complexity and comprehensive test suites that can be used for the assessment of state-of-the-art LLMs.